### v2

@@ Abstract
We consider the problem of approximating the regression function from noisy vector-valued data by an online learning algorithm using an appropriate reproducing kernel Hilbert space (RKHS) as prior. In an online algorithm, i.i.d. samples become available one by one by a random process and are successively processed to build approximations to the regression function. We are interested in the asymptotic performance of such online approximation algorithms and show that the expected squared error in the RKHS norm can be bounded by $C^{2}(m+1)^{-s/(2+s)}$, where $m$ is the current number of processed data, the parameter $0<s\leq 1$ expresses an additional smoothness assumption on the regression function, and the constant $C$ depends on the variance of the input noise, the smoothness of the regression function, and other parameters of the algorithm.
Keywords:
vector-valued kernel regression online algorithms convergence rates reproducing kernel Hilbert spaces
MSC:
65D15 65F08 65F1068W27
@@ 1 Introduction
In this paper, we consider the problem of learning the regression function from noisy vector-valued data using an appropriate RKHS as a prior. For relevant background on the theory of kernel methods, see Va1998 ; ScSm2002 ; CuSm2002 ; CuZh2007 ; StCh2008 and especially MiPo2005 ; CVT2006 ; CVTU2010 in the vector-valued case. Our focus is on obtaining estimates for the expectation of the squared error norm in the RKHS $H$ of approximations to the regression function which are built in an incremental way by so-called online algorithms. The setting we use is as follows: Let be given $N\leq\infty$ samples $(\omega_{m},y_{m})\in\Omega\times Y$, $m=0,\ldots,N-1$, of an input-output process $\omega\to y$, which are i.i.d. with respect to a (generally unknown) probability measure $\mu$ defined on $\Omega\times Y$. For simplicity, let $\Omega$ be a compact metric space, $Y$ a separable Hilbert space, and $\mu$ a Borel measure. What we are looking for is a regression function $f_{\mu}:\,\Omega\to Y$ which, in some sense, optimally represents the underlying input-output process. Algorithms for least-squares regression aim to find approximations to the solution
@@
[…]
for $f\in L^{2}_{\rho}(\Omega,Y)$ from the samples $(\omega_{m},y_{m}),m=0,\ldots,N-1$, where $\rho(\omega)$ is the marginal probability measure generated by $\mu(\omega,y)$ on $\Omega$ [^1] The symbol $\mathbb{E}$ denotes expectations of random variables with respect to the underlying probability space, which may vary from formula to formula but should be clear from the context. For the minimization problem (1) to be meaningful, one needs
[…]
@@
for $f\in L^{2}_{\rho}(\Omega,Y)$ is an ill-posed problem that makes no sense without further regularization, it is customary to add a prior assumption $f\in H$, where $H\subset L^{2}_{\rho}(\Omega,Y)$ is a set of functions $f:\;\Omega\to Y$ such that point evaluations $\omega\to f(\omega)$ are well-defined. Staying with the Hilbert space setting, candidates for $H$ are vector-valued RKHS
$H=\{f_{v}(\omega)=R^{\ast}_{\omega}v:\;\|f_{v}\|_{H}:=\|v\|_{V},\;v\in V\}$ (3)
introduced by means of a feature map $\mathbf{R}:=\{R_{\omega}\}_{\omega\in\Omega}$, i.e. by a family of bounded linear operators $R_{\omega}:\,Y\to V$ mapping into a separable Hilbert space $V$ called feature space. Detailed definitions and necessary assumptions on the feature map are given in the next section. Obviously, the RKHS $H$ and the feature space $V$ are isometric and can be identified.
The online algorithms considered in this paper build, starting from an initial guess $u^{(0)}\in V$, a sequence of successive approximations $u^{(m)}\in V$, where $u^{(m+1)}$ is a linear combination of the previous approximation $u^{(m)}$ and a term that includes the residual $y_{m}-R^{\ast}_{\omega_{m}}u^{(m)}$ with respect to the currently processed sample $(\omega_{m},y_{m})$. More precisely, the update formula in the feature space $V$ can be written in the form
$u^{(m+1)}(\omega)=\alpha_{m}(u^{(m)}+\mu_{m}R_{\omega_{m}}(y_{m}-R^{\ast}_{\omega_{m}}u^{(m)})),\qquad m=0,1,\ldots,N-1,$ (4)
@@
where $K(\omega,\theta):\,Y\to Y$, $\omega,\theta\in\Omega$, is the operator kernel of the RKHS determined by the feature map. The parameters $\alpha_{m},\mu_{m}$ are given by
@@
where the constants $11/2<t<$ 1 and $A>0$ will be properly fixed later. Following tradition, $\alpha_{m}$ and $\mu_{m}$ are called regularization and step-size parameters, respectively.
@@
of these iterations in the $V$ and $H$ norms, respectively. Here $u\in V$ is the unique minimizer of the minimization problem
[…]
These estimates hold under standard assumptions on the feature map, the parameters in (6) and the smoothness $s$ of $u\in V$ measured in a scale of smoothness spaces $V_{P_{\rho}}^{s}\subset V$ associated with the underlying covariance operator $P_{\rho}=\mathbb{E}(R_{\omega}R^{\ast}_{\omega})$. Our approach is an extension of earlier work GrOs2018 on Schwarz iterative methods in the noiseless case, where $y_{m}=f_{u}(\omega_{m})=R^{\ast}_{\omega_{m}}u$ was assumed.
The remainder of this paper is organized as follows: In Section 2 we introduce vector-valued RKHS, define $P_{\rho}$ and the associated scale of smoothness spaces $V_{P_{\rho}}^{s}$, and discuss properties of the minimization problem (7). This sets the stage for the analysis of our online learning algorithms in $V$ and allows us to formulate our main convergence result, namely Theorem 2.1. In Section 3 we review related results from the literature and compare them to our new result. In Section 4 we then provide the detailed proof of Theorem 2.1. In Section 5 we give further remarks on Theorem 2.1, show the divergence in expectation
@@
of the online algorithm (4)-(6) if (7) does not possess a unique minimizer $u\in V$ and consider a simple special case of ”learning” an element $u$ of a Hilbert space $V$ from noisy measurements of its coefficients with respect to a complete orthonormal system (CONS) in $V$.
@@ 2 Setting and main result
Let us first introduce our approach to vector-valued RKHS $H\subset L^{2}_{\rho}(\Omega,Y)$, where $\Omega$ is a compact metric space with Borel probability measure $\rho$ and $Y$ is a separable Hilbert space. Note that such a RKHS $H$ can be implicitly introduced by a family $\mathbf{R}=\{R_{\omega}\}_{\omega\in\Omega}$ of bounded linear operators $R_{\omega}:\,Y\to V$, where $V$ is another separable Hilbert space (we will tactitly assume that $V$ is infinite-dimensional). More precisely, we introduce the notation
@@
Under the simplifying conditions
$\|R^{\ast}_{\omega}v\|_{Y}=0\quad\rho-\mbox{a.e.}\quad\Longrightarrow\quad v=0\qquad\forall\;v\in V,$ (8)
and
$f_{v}(\omega)\in C(\Omega,Y)\qquad\forall\,v\in V,$ (9)
we then have $H\subset C(\Omega,Y)\subset L^{2}_{\rho}(\Omega,Y)$ and $f_{v}=0$ in $L^{2}_{\rho}(\Omega,Y)$ implies $v=0$. By definition, $H$ and $V$ are isometric, which allows us to easily switch between $H$ and $V$ in the sequel.
The condition (9) is the continuity of the operator family $\mathbf{R}$ in the strong operator topology and ensures Bochner integrability of functions from $\Omega$ into $Y$ and $V$, respectively, which appear in the formulas below. Because of the assumed compactness of $\Omega$, it also implies
@@
with some $\Lambda<\infty$. Another consequence is that the operator kernel
@@
associated with the vector-valued RKHS $H$ is a Mercer kernel. The condition (10) is equivalent to the uniform boundedness
of the operator kernel $K$. Furthermore, (10) is equivalent to the uniform boundedness of the operator family
@@
$\|P_{\omega}\|_{V}\leq\Lambda,\qquad\omega\in\Omega.$ (12)
For fixed $V$ and $\mathbf{R}$ satisfying the above properties, instead of solving the minimization problem (1) on $L^{2}_{\rho}(\Omega,Y)$, one now searches for the minimizer $u\in V$ of (7). The solution $u$ of this quadratic minimization problem on $V$, if it exists, must satisfy the necessary condition
@@
in $V$. Note that in general we cannot expect that $f_{u}=f_{\mu}$, since the closure $W$ of $H$ in $L^{2}_{\rho}(\Omega,Y)$ does not necessarily coincide with $L^{2}(\Omega,Y)$.
The operator $P_{\rho}:\,V\to V$ defined in (13), which plays the role of a covariance operator, is bounded and symmetric positive definite. Indeed,
[…]
and setting $w=v$ we also have $\mathrm{ker}(P_{\rho})=\{0\}$ since
$(P_{\rho}v,v)=\|f_{v}\|_{L^{2}_{\rho}(\Omega,V)}>0,\qquad 0\neq v\in V,$
by (8). The boundedness of $P_{\rho}:\,V\to V$, together with the estimate
@@
follows from (10). The spectrum of $P_{\rho}$ is thus contained in $[0,\Lambda]$. We also assume that $P_{\rho}$ is compact. A sufficient condition, which is often satisfied in applications, is the trace class property for $P_{\rho}$ which holds in particular if the operators $R_{\omega}$, $\omega\in\Omega$, have uniformly bounded finite rank.
Note that the assumptions on $\Omega$ and $\mathbf{R}$ can be weakened, see for example CVT2006 , and that the compactness of $P_{\rho}$ is used only as a technical simplification. In particular, the compactness assumption allows us to define the scale of smoothness spaces $V_{P_{\rho}}^{s}$, $s\in\mathbb{R}$, generated by $P_{\rho}$ using the complete orthonormal system (CONS) $\Psi:=\{\psi_{k}\}$ of eigenvectors of $P_{\rho}$ and associated eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\ldots>0$ with limit 00 in $V$ as follows: $V_{P_{\rho}}^{s}$ is the completion of $\mathrm{span}(\Psi)$ with respect to the norm
@@
which is well defined on $\mathrm{span}(\Psi)$ for any $s$. These spaces will appear in the investigation below.
For our convergence analysis in $V$, we now assume that
$\mathbb{E}(R_{\omega}y)\in\mathrm{ran}(P_{\rho}).$ (14)
In other words, we assume that (7) has a unique solution $u\in V$ for which (13) holds. For general $y$ with $\mathbb{E}(\|y\|_{Y}^{2})<\infty$, only $\mathbb{E}(R_{\omega}y)\in\mathrm{ran}(P^{1/2}_{\rho})$ can be established. It can also be proved that the functional $J(v)$ in (7) does not have a global minimum if (14) is violated (the infimum of $J(v)$ is then only approached if $v\to\infty$ in $V$ in a certain way). It will turn out that the condition (14) is necessary for the convergence in $V$ of the online method (4)-(6), see Subsection 5.3.
For a given prior RKHS $H$ induced by the operator family $\mathbf{R}$ with associated space $V$ and for given samples $(\omega_{m},y_{m})$, $m=0,\ldots,N-1$, with finite $N$, the standard regularization of the ill-posed problem (2) is to find the minimizer $u_{N}\in V$ of the minimization problem
@@
on $V$, where $\kappa_{N}>0$ is a suitable regularization parameter. Using the representer theorem for Mercer kernels MiPo2005 , this problem leads to a linear system with a typically dense and ill-conditioned $N\times N$ matrix. There is a huge body of literature, especially in the scalar-valued case $Y=\mathbb{R}$, devoted to setting up, analyzing and solving these problems for fixed $N$. Alternatively, one can restrict the minimization in (7) to bounded closed subsets in $V$ which, under our assumptions on $P_{\rho}$, are compact subsets of $L^{2}_{\rho}(\Omega,Y)$, see CuZh2007 .
We focus here on the online learning algorithm (4)-(6) for finding approximations to the minimizer $u\in V$ of (7) and are interested in the analysis of their asymptotic performance. For this purpose, we define the noise term
@@
which is a $Y$-valued random variable on $\Omega\times Y$ (to keep the notation short, the dependence of $\epsilon_{\omega}$ on $y$ is not explicitly shown). By (14) we have $\mathbb{E}(\epsilon_{\omega}|\omega)=f_{\mu}(\omega)-f_{u}(\omega)$ for any $\omega\in\Omega$. Also, the noise variance
@@
with respect to $f_{u}\in H$ is finite, since $\mathbb{E}(\|y\|_{Y}^{2})<\infty$ was assumed in the first place. The value of $\sigma_{H}$ depends on both the average size of the noise $y-f_{\mu}(\omega)$ on $\Omega$ measured in the $Y$ norm and the $L_{\rho}^{2}(\Omega,Y)$ distance of $f_{\mu}$ from $W$.
The online algorithm (4)-(6) is a particular instance of a randomized Schwarz approximation method associated with $\mathbf{R}$. Its noiseless version, where $y_{m}=R_{\omega_{m}}^{\ast}u$, was studied in GrOs2018 under the assumption $u\in V^{s}_{P_{\rho}}$, $0\leq s\leq 1$, where $\alpha_{m}$ was as in (6) but $\mu_{m}$ was determined by a steepest descent rule. Our goal in this paper is to derive convergence results for the expected squared error $\mathbb{E}(\|u-u^{(m)}\|^{2}_{V})$, $m=1,2,\ldots$. As expected, such estimates again require additional smoothness assumptions on $u$ in the form $u\in V_{P_{\rho}}^{s}$ with $0<s\leq 1$. However, in contrast to the noiseless case GrOs2018 , they include a dependence on the noise variance $\sigma_{H}^{2}$ in addition to the dependence on the initial error and the smoothness of $u$. The prize of convergence is a certain decay of the step-sizes $\mu_{m}\to 0$, as assumed in (6), which is typical for stochastic approximation algorithms. Our main result is as follows:

### v5

@@ Abstract
We consider the problem of approximating the regression function $f_{\mu}:\,\Omega\to Y$ from noisy $\mu$-distributed vector-valued data $(\omega_{m},y_{m})\in\Omega\times Y$ by an online learning algorithm using a reproducing kernel Hilbert space $H$ (RKHS) as prior. In an online algorithm, i.i.d. samples become available one by one via a random process and are successively processed to build approximations to the regression function. Assuming that the regression function essentially belongs to $H$ (soft learning scenario), we provide estimates for the expected squared error in the RKHS norm of the approximations $f^{(m)}\in H$ obtained by a standard regularized online approximation algorithm. In particular, we show an order-optimal estimate
$\mathbb{E}(\|\epsilon^{(m)}\|_{H}^{2})\leq C(m+1)^{-s/(2+s)},\qquad m=1,2,\ldots,$
where $\epsilon^{(m)}$ denotes the error term after $m$ processed data, the parameter $0<s\leq 1$ expresses an additional smoothness assumption on the regression function, and the constant $C$ depends on the variance of the input noise, the smoothness of the regression function, and other parameters of the algorithm. The proof, which is inspired by results on Schwarz iterative methods GrOs2018 in the noiseless case, uses only elementary Hilbert space techniques and minimal assumptions on the noise, the feature map that defines $H$ and the associated covariance operator.
@@ 1 Introduction
In this paper, we consider the problem of learning the regression function from noisy vector-valued data using an appropriate RKHS as a prior. For relevant background on the theory of kernel methods, see Va1998 ; ScSm2002 ; CuSm2002 ; CuZh2007 ; StCh2008 and especially MiPo2005 ; CVT2006 ; CVTU2010 in the vector-valued case. Our focus is to obtain estimates for the expectation of the squared error norm in the RKHS $H$ of approximations to the regression function. These approximations are constructed in an incremental way by so-called online algorithms. The setting we use is as follows: Let be given $N\leq\infty$ samples $(\omega_{m},y_{m})\in\Omega\times Y$, $m=0,\ldots,N-1$, of an input-output process $\omega\to y$, which are i.i.d. with respect to a (generally unknown) probability measure $\mu$ defined on $\Omega\times Y$. Let $\Omega$ be a compact metric space, $Y$ be a separable Hilbert space, and $\mu$ be a Borel measure. We are looking for a regression function $f_{\mu}:\,\Omega\to Y$ that optimally represents the underlying input-output process in some sense. Algorithms for least-squares regression aim to find approximations to the solution
@@
$\mathbb{E}(\|f(\omega)-y\|_{Y}^{2})=\int_{\Omega\times Y}\|f(\omega)-y\|_{Y}^{2}\,d\mu(\omega,y)\longmapsto\min$ (1)
for $f\in L^{2}_{\rho}(\Omega,Y)$ from the samples $(\omega_{m},y_{m}),m=0,\ldots,N-1$, where $\rho(\omega)$ is the marginal probability measure generated by $\mu(\omega,y)$ on $\Omega$. [^1] The symbol $\mathbb{E}$ denotes expectations of random variables with respect to the underlying probability space, which may vary from formula to formula but should be clear from the context. For the minimization problem (1) to be meaningful, one needs
$\mathbb{E}(\|y\|_{Y}^{2})=\int_{\Omega\times Y}\|y\|_{Y}^{2}\,d\mu(\omega,y)=\int_{\Omega}\mathbb{E}(\|y\|_{Y}^{2}|\omega)\,d\rho(\omega)<\infty.$
@@
for $f\in L^{2}_{\rho}(\Omega,Y)$ is an ill-posed problem that makes no sense without further regularization, it is customary to add a prior assumption $f\in H$, where $H\subset L^{2}_{\rho}(\Omega,Y)$ is a set of functions $f:\;\Omega\to Y$ such that point evaluations $\omega\to f(\omega)$ are continuous maps. Staying with the Hilbert space setting, candidates for $H$ are vector-valued RKHS
$H=\{f_{v}(\omega)=R^{\ast}_{\omega}v:\;\|f_{v}\|_{H}:=\|v\|_{V},\;v\in V\}$
introduced by means of a family $\mathbf{R}:=\{R_{\omega}\}_{\omega\in\Omega}$ of bounded linear operators $R_{\omega}:\,Y\to V$ mapping into a separable Hilbert space $V$ called feature space. Detailed definitions and necessary assumptions on the feature map $\mathbf{R}$ are given in the next section. [^2] In the literature MiPo2005 ,(StCh2008, , Chapter 4), feature maps are typically introduced as operator families acting from $V$ to $Y$, which, in our case, would be the family $\mathbf{R}^{\ast}$ of Hilbert adjoints $R_{\omega}^{\ast}:\,V\to Y$. For the purpose of this paper, we stay close to the notation used in GrOs2018 and call $\mathbf{R}$ feature map, see also CVT2006 . Obviously, if $f_{v}(\omega)=0$ for all $\omega\in\Omega$ implies $v=0$ then $\mathbf{R}^{\ast}$ is an isometry and the RKHS $H$ and the feature space $V$ can be identified.
The online algorithms considered in this paper start from an initial guess $u^{(0)}\in V$ and build a sequence of successive approximations $u^{(m)}\in V$, where $u^{(m+1)}$ is a linear combination of the previous approximation $u^{(m)}$ and a term that includes the residual $y_{m}-R^{\ast}_{\omega_{m}}u^{(m)}$ with respect to the currently processed sample $(\omega_{m},y_{m})$. More precisely, the update formula in the feature space $V$ can be written in the form
$u^{(m+1)}(\omega)=\alpha_{m}(u^{(m)}+\mu_{m}R_{\omega_{m}}(y_{m}-R^{\ast}_{\omega_{m}}u^{(m)})),\qquad m=0,1,\ldots,N-1.$ (3)
@@
where $K(\omega,\theta):\,Y\to Y$, $\omega,\theta\in\Omega$, is the operator kernel of the RKHS determined by the feature map.  [^3] The formula (4) shows that, in order to execute the algorithm, the explicit use of $V$ and the feature map $\mathbf{R}$ can be avoided if the operator kernel $K$ and not the feature map is given, which is often the case. In the convergence proofs, it is more convenient to use (3). The parameters $\alpha_{m},\mu_{m}$ are specifically given by
@@
where the constants $1/2<t<1$ and $A>0$ will be properly fixed later. Following tradition, $\alpha_{m}$ and $\mu_{m}$ are called regularization and step-size parameters, respectively.
@@
of these iterations in the $V$ and $H$ norms, respectively. Here $u\in V$ is the unique minimizer of the minimization problem [^4] The existence of such a $u$ is a strong assumption necessary for investigating convergence in the RKHS norm and is called soft learning scenario in the literature, see, e.g., FiSt2020 . Our method of proof does not automatically extend to the hard learning scenario, where only $f_{\mu}\in L^{2}_{\rho}(\Omega,Y)$ is assumed and $L^{2}_{\rho}(\Omega,Y)$ convergence is studied.
$J(v):=\mathbb{E}(\|f_{v}-y\|_{Y}^{2})=\int_{\Omega\times Y}\|R^{\ast}_{\omega}v-y\|_{Y}^{2}\,d\mu(\omega,y)\longmapsto\min.$ (6)
These estimates hold under standard assumptions on the feature map $\mathbf{R}$ and the operator kernel $K$, the parameters in (5) and the smoothness $s$ of $u\in V$ measured in a scale of smoothness spaces $V_{P_{\rho}}^{s}\subset V$ associated with the underlying covariance operator $P_{\rho}=\mathbb{E}(R_{\omega}R^{\ast}_{\omega})$.
Our approach is an extension of earlier work GrOs2018 on Schwarz iterative methods in the noiseless case, where $y_{m}=f_{u}(\omega_{m})=R^{\ast}_{\omega_{m}}u$ was assumed. In our opinion, the main contribution of this paper is providing a proof of a generic convergence result for online learning in the general vector-valued case which uses only basic Hilbert space techniques. In particular, compactness assumptions on the covariance operator $P_{\rho}$ do not play a role other than to simplify the presentation of proofs, and structural assumptions on the noise $y-f_{\mu}(\omega)$ are not needed. We note that the online learning version studied in this paper can be reinterpreted as the iterative solution of an inverse problem for the equation $R_{\omega}^{\ast}u=y$ (equality in $L^{2}_{\rho}(\Omega,Y)$) with random right-hand side $y$. Iterative regularization of inverse problems with deterministic noise is studied in (EHN2000, , Chapter 6) using similar tools. The connection between kernel regression and statistical inverse problems is well-known, see VRCGO2005 , but we are not aware of a result similar to our Theorem 2.1 covering the online case. More sophisticated tools (for instance, concentration inequalities valid under certain assumptions on the noise, integral operator techniques, taking into account decay properties of the spectrum of $P_{\rho}$) lead to stronger results that apply to the hard learning scenario as well. Furthermore, we do not claim relevance for particular applications. Motivation for the vector-valued case comes from multitask learning (here, $Y=\mathbb{R}^{d}$) and functional learning, see the references in ARL2012 ; KDPCRA2015 ; LMMG2024 . In most of the current applications, constructions involving $\mathbb{R}$-valued kernels are used to cover the vector-valued case. In Section 3, we provide an example with $Y=H^{1}_{0}(G)$ following KDPCRA2015 .
The remainder of this paper is organized as follows: In Section 2 we introduce vector-valued RKHS in terms of feature maps $\mathbf{R}$ and spaces $V$, define $P_{\rho}$ and the associated scale of smoothness spaces $V_{P_{\rho}}^{s}$, and discuss properties of the minimization problem (6). This sets the stage for the analysis of our online learning algorithm in $V$ and allows us to formulate our main convergence result, namely Theorem 2.1. In Section 3 we review related results from the literature. In Section 4 we then provide the detailed proof of Theorem 2.1. In Section 5 we give further remarks on Theorem 2.1, show the divergence in expectation
@@
of the online algorithm (3)-(5) if (6) does not possess a unique minimizer $u\in V$, and consider a simple special case of ”learning” an element $u$ of a Hilbert space $V$ from noisy measurements of its coefficients with respect to a complete orthonormal system (CONS) in $V$.
@@ 2 Setting and main result
Let us first introduce our approach to vector-valued RKHS $H\subset L^{2}_{\rho}(\Omega,Y)$, where $\Omega$ is a compact metric space with Borel probability measure $\rho$ and $Y$ is a separable Hilbert space. [^5] The facts stated below without proof can be found in CVT2006 ; PaRa2016 . The assumptions on $\Omega$ can be weakened. In the learning problem under consideration, $\rho$ is the marginal measure induced from a Borel probability measure $\mu$ on $\Omega\times Y$. Such an RKHS $H$ can be implicitly introduced by a family $\mathbf{R}=\{R_{\omega}\}_{\omega\in\Omega}$ of bounded linear operators $R_{\omega}:\,Y\to V$, called feature map, where $V$ is another separable Hilbert space (we will tacitly assume that $V$ is infinite-dimensional). More precisely, we introduce the notation
@@
Without loss of generality, we assume that $f_{v}(\omega)=0$ for all $\omega\in\Omega$ implies $v=0$ (otherwise, replace $V$ by its subspace $(\cap_{\omega\in\Omega}\mathrm{ker}(R_{\omega}^{\ast}))^{\perp}$). That is, the RKHS $H$ and the space $V$ can be identified, which allows us to easily switch between $H$ and $V$ in the sequel.
We impose the following conditions on the feature map $\mathbf{R}$: We assume uniform boundedness
@@
with some $\Lambda<\infty$ and (weak) measurability, i.e., the scalar function $(R_{\omega}y,v)_{V}$ is Borel measurable on $\Omega$ for each pair $(y,v)\in Y\times V$. By (7), the functions $f_{v}\in H$ are bounded on $\Omega$ which, together with the measurability assumption, implies $H\subset L^{2}(\Omega,Y)$ for any Borel probability measure $\rho$ on $\Omega$. [^6] For simplicity, we silently identify $f_{v}$ with its equivalence class, whenever this is formally necessary.
The condition (7) is equivalent to the uniform boundedness
of the operator kernel
@@
associated with the vector-valued RKHS $H$. Furthermore, (7) is equivalent to the uniform boundedness of the operator family
@@
$\|P_{\omega}\|_{V}\leq\Lambda,\qquad\omega\in\Omega.$
Moreover, weak measurability of the operator families $K(\omega,\theta)$, $R^{\ast}_{\omega}$ and $P_{\omega}$ on $\Omega\times\Omega$ and $\Omega$, respectively, follows then from the weak measurability of $\mathbf{R}$ as well. This ensures that Bochner integrals of functions from $\Omega$ into $Y$ and $V$, respectively, which appear in the formulas below, are well-defined.
For fixed $V$ and $\mathbf{R}$ satisfying the above properties, instead of solving the minimization problem (1) on $L^{2}_{\rho}(\Omega,Y)$, one now searches for the minimizer $u\in V$ of (6). The solution $u$ to this quadratic minimization problem on $V$, if it exists, must satisfy the necessary condition
@@
in $V$. Note that in general we cannot expect that $f_{u}=f_{\mu}$, since the closure $W$ of $H$ in $L^{2}_{\rho}(\Omega,Y)$ does not necessarily coincide with $L^{2}_{\rho}(\Omega,Y)$. For our convergence analysis in $V$, we require that
$\mathbb{E}(R_{\omega}y)\in\mathrm{ran}(P_{\rho}).$ (10)
In other words, we assume that (6) has a unique solution $u\in V$ for which (9) holds. For general $y$ with $\mathbb{E}(\|y\|_{Y}^{2})<\infty$, only $\mathbb{E}(R_{\omega}y)\in\mathrm{ran}(P^{1/2}_{\rho})$ can be established. It can also be proven that the functional $J(v)$ in (6) does not have a global minimum if (10) is violated (the infimum of $J(v)$ is then only approached if $v\to\infty$ in $V$ in a certain way). It will turn out that the condition (10) is necessary for the convergence in $V$ of the online method (3)-(5), see Subsection 5.3.
The operator $P_{\rho}:\,V\to V$ in (9), which plays the role of a covariance operator, is bounded and symmetric positive definite. Indeed, by definition we have
$(P_{\rho}v,w)_{V}$ $:=$ $\int_{\Omega\times Y}(P_{\omega}v,w)_{V}\,d\mu(\omega,y)=\int_{\Omega}(P_{\omega}v,w)_{V}\,d\rho(\omega)$
$=$ $\int_{\Omega}(R^{\ast}_{\omega}v,R^{\ast}_{\omega}w)_{Y}\,d\rho(\omega)=(f_{v},f_{w})_{L^{2}_{\rho}(\Omega,Y)}=(v,P_{\rho}w)_{V},\quad v,w\in V,$
and
$(P_{\rho}v,v)=\|f_{v}\|_{L^{2}_{\rho}(\Omega,Y)}^{2}\geq 0,\qquad v\in V.$
Even though it may happen that $P_{\rho}$ is not injective for certain $\rho$, we will from now on assume that $\mathrm{ker}(P_{\rho})=\{0\}$ (otherwise, replace $V$ by its subspace $\{v\in V:\;\|f_{v}\|_{L^{2}_{\rho}(\Omega,Y)}=0\}^{\perp}$). The boundedness of $P_{\rho}:\,V\to V$, together with the estimate
@@
follows from (7). The spectrum of $P_{\rho}$ is thus contained in $0,\Lambda]$.
In the formulation of the results and proofs below, we will additionally assume that $P_{\rho}$ is compact, which is the case in all known applications. A sufficient condition for compactness is the trace class property for $P_{\rho}$, which holds in particular if the operators $R_{\omega}$, $\omega\in\Omega$, have uniformly bounded finite rank. In the scalar-valued case $Y=\mathbb{R}$, the trace class property holds if the associated kernel function $k:\,\Omega\times\Omega\to\mathbb{R}$ satisfies
$\int_{\Omega}k(\omega,\omega)\,d\rho(\omega)<\infty,$
which is automatically satisfied if the boundedness condition (8) holds. The compactness assumption enables us to define the scale of smoothness spaces $V_{P_{\rho}}^{s}$, $s\in\mathbb{R}$, generated by $P_{\rho}$ by simply using the complete orthonormal system (CONS) $\Psi:=\{\psi_{k}\}$ of eigenvectors of $P_{\rho}$ and associated eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\ldots>0$ with limit 0 in $V$ as follows: $V_{P_{\rho}}^{s}$ is the completion of $\mathrm{span}(\Psi)$ with respect to the norm
@@
which is well defined on $\mathrm{span}(\Psi)$ for any $s$. These spaces will appear in the investigation below. For a noncompact $P_{\rho}$, the norm of the spaces $V_{P_{\rho}}^{s}$ can be defined using functional calculus, i.e., by replacing series representations using eigenfunction expansions by integrals with respect to the underlying spectral family of the operator $P_{\rho}$. For example, this is worked out in EHN2000 in connection with iterative regularization methods of inverse problems (for $s>0$, the space $V_{P_{\rho}}^{s}$ equals ${X}_{s/2}$ defined in (EHN2000, , (3.29)) if we replace $T^{\ast}T$ by $P_{\rho}$).
For a given prior RKHS $H$ induced by the feature map $\mathbf{R}$ with associated feature space $V$ and for given samples $(\omega_{m},y_{m})$, $m=0,\ldots,N-1$, with finite $N$, the standard regularization of the ill-posed problem (2) is to find the minimizer $u_{N}\in V$ of the minimization problem
@@
on $V$, where $\kappa_{N}>0$ is a suitable regularization parameter. Using the representer theorem for Mercer kernels MiPo2005 ; CVT2006 , this problem leads to a linear system with a typically dense and ill-conditioned $N\times N$ matrix. There is a huge body of literature, especially in the scalar-valued case $Y=\mathbb{R}$, devoted to setting up, analyzing and solving these problems for fixed $N$. Alternatively, one can restrict the minimization in (6) to bounded, closed subsets in $V$ which, under our assumptions on $P_{\rho}$, are compact subsets of $L^{2}_{\rho}(\Omega,Y)$, see CuZh2007 .
Here, we focus on the online learning algorithm (3)-(5) for finding approximations to the minimizer $u\in V$ of (6) and are interested in the analysis of their asymptotic performance. For this, we define the noise term
@@
which is a $\mu$-distributed $Y$-valued random variable on $\Omega\times Y$ (to keep the notation short, the dependence of $\epsilon_{\omega}$ on $y$ is not explicitly shown). By (10) we have $\mathbb{E}(\epsilon_{\omega}|\omega)=f_{\mu}(\omega)-f_{u}(\omega)$ for any $\omega\in\Omega$. Also, the noise variance
@@
with respect to $f_{u}\in H$ is finite, since $\mathbb{E}(\|y\|_{Y}^{2})<\infty$ was assumed in the first place. The value of $\sigma_{H}$ depends on both the average size of the noise $y-f_{\mu}(\omega)$ on $\Omega$ measured in the squared $Y$ norm and the $L_{\rho}^{2}(\Omega,Y)$ distance of $f_{\mu}$ from $W$.
The online algorithm (3)-(5) is a particular instance of a randomized Schwarz iteration method associated with $\mathbf{R}$. Its noiseless version, where $y_{m}=R_{\omega_{m}}^{\ast}u$, was studied in GrOs2018 under the assumption that $u\in V^{s}_{P_{\rho}}$, $0\leq s\leq 1$, where $\alpha_{m}$ was as in (5) but $\mu_{m}$ was determined by a steepest descent rule. Our goal in this paper is to derive convergence results for the expected squared error $\mathbb{E}(\|u-u^{(m)}\|^{2}_{V})=\mathbb{E}(\|f_{u}-f^{(m)}\|^{2}_{H})$, $m=1,2,\ldots$. As expected, such estimates again require additional smoothness assumptions on $u$ in the form $u\in V_{P_{\rho}}^{s}$ with $0<s\leq 1$. However, unlike the noiseless case GrOs2018 , these estimates also depend on the noise variance $\sigma_{H}^{2}$ in addition to the dependence on the initial error and the smoothness of $u$. The price of convergence is a certain decay of the step-sizes $\mu_{m}\to 0$, as assumed in (5), which is typical of stochastic approximation algorithms. Our main result is as follows: